Write an equation in slope-intercept form of the line satisfying the given conditions. Perpendicular to the line y = 2x + 3; containing the point (-2, -4).
Answer by dfrazzetto(283)


slope of original line m=2
perpendicular slope is inverse reciprocal m= -1/2
y = (-1/2)x + b
-4 = (-1/2)(-2) + b
-4 = 1 + b
b = -5



y = (-1/2)x - 5